what is the slope for y= x/5 + 4/5
Found 2 solutions by thepraveenrakki, lenny460:
Answer by thepraveenrakki(3)

y = x/5 + 4/5
y = 1/5x + 4/5
(According to the general equation, y = mx + c, the slope of the graph/equation is 1/5.)
Thus, the slope is 1/5.

Answer by lenny460(1073)

what is the slope for y= x/5 + 4/5

The Slope Intercept Formula:
y = mx + b
Therefore:
y = x/5 + 4/5 or
y = 1x/5 + 4/5
Therefore the Slope is:
1/5



Lennox Obuong
Algebra Student
Email: obuong3@aol.com